Circular No. 7864

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002bv IN UGC 4042 P. Martin and W. D. Li, University of California at Berkeley (UCB),

report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova (mag ∼ 17.8) on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Mar. 29.2 and 30.2 UT. SN 2002bv is located at α = 7^{h}50^{m}01^{s}.13, δ = +30^{o}01′32′′.2 (equinox 2000.0), which is 14′′.3 east and 2′′.3 north of the nucleus of UGC 4042. A KAIT image taken on Jan. 31.2 showed nothing at this position (limiting mag ∼ 19.5).

SUPERNOVA 2002bw IN PGC 59990 Li also reports the LOTOSS discovery of an apparent supernova (mag

∼ 17.4) on unfiltered KAIT images taken on Mar. 30.5 and Apr. 1.5 UT. SN 2002bw is located at α = 17^{h}16^{m}35^{s}.60, δ = −10^{o}20′31′′.8 (equinox 2000.0), which is 3′′.1 west and 7′′.3 north of the nucleus of PGC 59990. A KAIT image taken on 2001 Sept. 9.3 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002bx IN IC 2461 M. Ganeshalingam and Li report the LOTOSS discovery of another

apparent supernova on unfiltered KAIT images taken on Mar. 31.2 (mag ∼ 16.9) and Apr. 1.2 UT (mag ∼ 16.8). SN 2002bx is located at α = 9^{h}19^{m}54^{s}.98, δ = +37^{o}12′08′′.3 (equinox 2000.0), which is 36′′.6 west and 40′′.6 north of the nucleus of IC 2461. A KAIT image of the same field on Mar. 11.2 showed nothing at this position (limiting mag ∼ 18.5). An independent discovery of SN 2002bx at mag 16.2 was made by T. Boles, Coddenham, England, on images obtained on Apr. 1.943 in the course of the U.K. Supernova Patrol; Boles provides position end figures 55^{s}.01, 08′′.9.

SUPERNOVA 2002bu IN NGC 4242 K. Ayani and T. Kawabata, Bisei Astronomical Observatory (BAO);

and H. Yamaoka, Kyushu University, report that a low-resolution spectrum (range 400–700 nm; resolution 0.7 nm) of SN 2002bu (IAUC 7863), taken with the BAO 1.01-m reflector on Mar. 31.66 UT, shows a flat continuum with strong and narrow (FWHM ∼ 1100 km/s) Balmer emission lines, indicating that it is of type IIn. The redshifts of emission lines are ∼ 600 km/s, which is consistent with that of NGC 4242.

2002 April 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green